Grade 11Physical Chemistry

Q.200 gm of marble chips are dropped into 1 kg of HCl sol containing one-tenth of its weight of the pure acid.How much of chips will remain undissolved.What weight of anhydrous calcium chloride and CO2 gas could be obtained?

SF
9 Years agoGrade 11

1 Answer

Samridhi
9 Years ago
Given: mass of calcium carbonate (CaCO3) = 200 gHydrochloric acid (HCl)= 100 * 1/ 10 ( one tenth of 1000 gm is 100 gm is pure)=100gReaction:CaCO3 + 2HCl = CaCl2 +CO2 +H2O molar masses of compounds in reaction ;CaCO3 = 100g2HCl = 142 gCaCl2 = 111g CO2 = 44gNow ...142g of HCl reacts with 100g of CaCO3 Then ...100 g of HCl will react with CaCO3 = 100*100/142 = 70.42gBut we have chips ie CaCO3 =200 gChips left are 200-70.42= 129.58142g of HCl produce 111f of CaCl2 100g of HCl will produce CaCl2 = 111*100/142 = 78.17g142g of HCl produces 44g of CO2100g of HCl will produce CO2 = 44*100/142 = 30.98g